OpenStudy (anonymous):

[(√75a) + (√12a) - (√27a)]/(√3a)

OpenStudy (shubhamsrg):

sqrt(a)/sqrt(b) = sqrt(a/b) try now

OpenStudy (anonymous):

what? o.O

OpenStudy (shubhamsrg):

well (a+b)/c = a/c + b/c use this thing,what do you get?

OpenStudy (anonymous):

i dont know...

OpenStudy (shubhamsrg):

hmmn..you want the ans only na ?

OpenStudy (anonymous):

i pretty much can't do anything with square roots...they confuse me

OpenStudy (shubhamsrg):

try and simplify..for eg,,you can write sqrt(75) as sqrt(5*5*3) =5sqrt(3) can you simplify rest now ?

OpenStudy (anonymous):

nope...

OpenStudy (shubhamsrg):

hmmn,,ans is 4..:|

OpenStudy (anonymous):

thanks >.< im horrible at this